OpenStudy (anonymous):

Write a pseudocode algorithm to compute the product of the first n positive integers. How many multiplications does your algorithm perform?

OpenStudy (anonymous):

we could do a loop... x=0 for i = 1 to n x=x*i next I guess the algorithm does this with 1 multiplication, but perhaps it could be said to be done with n multiplications. I think then if n = 5 then we get the product of 1 * 1 = 1 1 * 2 = 2 2 * 3 = 6 6 * 4 = 24 24 * 5 = 120 1 * 2 * 3 * 4 * 5 = 120 there might be another way with a geometric or arithmetic sequence though

OpenStudy (anonymous):

Thanks for the answer... I waited and grew impatient. I came up with a similar algorithm and also stated that there were n multiplications. Since both of us came up with this, it'll be right :)